The shareholders of Flannery Company have voted in favor of a buyout offer from Stultz Corporation. Information about each firm is given here:

Flannery Stultz
  Price–earnings ratio 12 36
  Shares outstanding 90,000 360,000
  Earnings $ 220,000 $ 880,000

Flannery’s shareholders will receive one share of Stultz stock for every three shares they hold in Flannery.

a-1.

What will the EPS of Stultz be after the merger? (Do not round intermediate calculations and round your answer to 3 decimal places, e.g., 32.161.)

  EPS $

a-2.

What will the PE ratio be if the NPV of the acquisition is zero? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

  PE ratio

  

b.

What must Stultz feel is the value of the synergy between these two firms? (Do not round intermediate calculations.)

  

  Synergy value $
Answer #1

a-1). EPS = Total Earnings / Total number of shares outstanding

= [Fannery's Earnings + Stultz's Earnings] / [{(1/3) * Flannery's Shares} + Stultz's Shares]

= [$220,000 + $880,000] / [{(1/3) * 90,000} + 360,000]

= $1,100,000 / 390,000 = $2.821

a-2). The market price of Stultz will remain unchanged if it is a zero NPV acquisition.

P = [P.E. Ratio * Stultz's Earnings] / Stultz's Shares

= [36 * $880,000] / 360,000 = $31,680,000 / 360,000 = $88

If the acquisition has a zero NPV, the stock price should remain unchanged. Therefore, the new PE will be:

PE = $88 / $2.821 = 31.2

b). The value of Flannery to Stultz must be the market value of the company since the NPV of the acquisition is zero

V = Earnings * PE Ratio

= $220,000 * 12 = $2,640,000

Cost = Number of shares offered * Share Price

= [(1/3)*90,000] * $88 = $2,640,000

NPV = 0 = V + \Delta V - Cost

0 = $2,640,000 + \Delta V - $2,640,000

\DeltaV = $0